### v5

# Intensity-Efficient Allocations

Abstract
This paper proposes a refinement of Pareto-efficient allocations for situations where, in addition to having ordinal preferences, agents also have ordinal intensities: they can make comparisons such as “I prefer $a$ to $b$ more than I prefer $c$ to $d$, without necessarily being able to quantify them. A rank-based criterion for interpersonal comparisons of such ordinal intensities is introduced for this new analytical environment. Building on this, an allocation is defined to be intensity-efficient if it is Pareto-efficient with respect to the agents’ preferences and also such that when another Pareto-efficient allocation assigns the same pairs of items to the same pairs of agents but in a “flipped” way, the former allocation assigns the commonly preferred item in every such pair to the agent who prefers it more. Conditions are established under which such Pareto-refining allocations exist. The potential usefulness of this theory in matching problems is illustrated with a quadratic-time extension of the Random Priority (RP) algorithm that returns an allocation which intensity-dominates RP’s Pareto-efficient one.
“Suppose I am left with a ticket to a Mozart concert I am unable to attend and decide to give it to one of my closest friends. Which friend should I actually give it to? One thing I will surely consider in deciding this is which friend of mine would enjoy the concert most.”
John C. Harsanyi (1987)
“The problem I have with utilitarianism is not that it is excessively rational, but that the epistemological foundations are weak. My problem is: What are those objects we are adding up? I have no objection to adding them up if there’s something to add.”
Kenneth J. Arrow (1987)
1 Introduction
Distributional principles that are used in the assignment of goods to individuals can be evaluated jointly on the realism of their assumptions and the appeal—according to some justice criterion—of their prescribed allocations. Arguably the most important such economic principle is Pareto-efficiency: assuming only information on agents’ preference rankings over the relevant goods, its notion of optimality is defined by the requirement that no changes can make some individuals better off without hurting others. Yet it is well-known that among the typically many Pareto-efficient allocations there are often some which violate intuitive notions of distributive justice. Appreciation of these facts raises the question of whether one can narrow down and improve upon Pareto’s criterion by considering alternative principles which, in addition to agents’ preferences, also incorporate information on their preference intensities.
Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities. Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered. Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion formalized in Harsanyi (1955) is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Roemer, 1996; Moulin, 2004).
Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Dietrich, 2025; Samuelson, 2025). More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982). This stands in contrast to some contemporary views in the field of cognitive psychology and decision-making, according to which “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” (Vlaev et al., 2011, p. 551). To the extent that comparing options is easier and more reliable than assigning values to them, as the above survey and several studies cited therein suggest, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made. Specifically, instead of cardinal-welfarist or Arrowian profiles that, respectively consist of interpersonally comparable cardinal utility functions—possibly not arising from expected-utility preferences (Nebel, 2024; Weymark, 2005; Harvey, 1999)—or preference orderings, this domain comprises profiles of generally ordinal preference intensity orderings, i.e. collections of quaternary relations that reflect the agents’ preference intensity comparisons. This analytical choice assumes (and this assumption is validated in Section 5, Proposition 3) that, similar to the way in which preferences are elicited, e.g. in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
This new and general framework where both preference and preference-intensity information is available to the social planner naturally raises the question of how such information might be used—as per the goals that we set out above—to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ preference intensities and mitigate the multiplicity of Pareto-efficient allocations. Similar to utilitarianism and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons. Unlike those notions, however, where such comparisons are based on agents’ utilities (Baccelli, 2023; Echenique et al., 2023; d’Aspremont, 2018; Fleurbaey and Hammond, 2004), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful. In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more. In Section 3 we formalize and defend this Interpersonal Comparisons of Ordinal Intensities new postulate/informational basis for social choice and welfare. We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (Baccelli, 2023; d’Aspremont, 2018).
In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just. Specifically, we define an allocation $x$ to be intensity-efficient if it is Pareto-efficient and also such that, whenever another Pareto-efficient allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more. Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile, without further restrictions, when there are three agents and alternatives. Yet we also show that with four or more existence is not guaranteed because the underlying dominance relation may be cyclic without further assumptions. For the case of $n$ agents and items we then establish a sufficient condition which requires profiles to be (effectively) homeo-monotonic and whose key part could be summarized informally thus: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any two consecutively-ranked items in this subset, then this ordering cannot be violated in how much they prefer any non-consecutively ranked items. Furthermore, although this condition is not necessary for existence, we prove that a weaker version of it is indeed satisfied by every profile that has an intensity-efficient allocation.
Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible. Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide, although they do so under an intuitive additional condition that we introduce. The potential discrepancy between these two intensity-incorporating welfare concepts that may arise in such cardinal environments is itself of potential usefulness in applications, because either notion can be used as a secondary selection criterion in those situations where the other one predicts multiple allocations.
In Section 5, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon Pareto-efficiency in the direction of intensity-efficiency. We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited efficiently (specifically, in quadratic time) by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class we study. Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) one and produces—in quadratic time in the number of items/agents—an allocation that intensity-dominates the Pareto-efficient allocation that is found by RP. While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm. We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and which focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
1.1 Related Literature
This paper’s contribution lies at the intersection of welfare economics, social choice and matching theory. Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Kaneko and Nakamura, 1979; Roemer, 1996; Moulin, 2004), Pareto-efficient egalitarian-equivalence (Pazner and Schmeidler, 1978) or relative egalitarianism (Kalai and Smorodinsky, 1975; Imai, 1983; Sprumont, 2013), the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process. Unlike these refinements, the hereby proposed one is the first that is operational in the general class of ordinal preference-intensity profiles, and builds on a novel interpersonal-comparability criterion that is distinct from interpersonal comparisons of utility differences, which the above notions rely on. Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain. The latter problem is important and deserves separate treatment.
Intensity-efficient allocations and their domain of application are also distinct from those dictated by the Borda scoring rule (Borda, 1781; Young, 1974; Maskin, 2025). Although this rule is defined by a particular normalization of agents’ ordinal utility functions over alternatives, in practice it is often used with a cardinal interpretation due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where every agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart. The class of ordinal intensity relations that we consider does not impose this restriction. Yet at the same time, for the main results of this paper we assume a strictness condition on agents’ intensities which rules out such “linear” intensity profiles.
In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs in the allocation process toward maximizing utilitarian welfare. Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer. While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
2 Notation, Preliminaries and Measurement-Theoretic Foundations
By $A:=\{a_{1},\ldots,a_{n}\}$ and $V:=\{v_{1},\ldots,v_{n}\}$ we denote, respectively, the finite sets of choice alternatives and agents. To simplify the assignment problem under study we assume that these sets have equal sizes. When subscripts are unnecessary we also write $a,b\in A$. Similarly, when no confusion arises we will refer to agent $v_{l}$ by their subscript, $l$. The preference intensity relation of agent $l\leq n$ on $A$ is denoted by $\dot{\succsim}_{l}$, with $\dot{\succ}_{l}$ and $\dot{\sim}_{l}$ its asymmetric and symmetric parts. In line with extensive-measurement theory (Krantz et al., 1971; Pfanzagl, 1971; Roberts, 1979), these are quaternary relations on $A$ or, equivalently, binary relations on $A\times A$. The statement $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ will be interpreted as “$a$ is preferred to $b$ at least as much as $c$ is to dd” when the first option in each pair is the (possibly weakly) preferred one at that pair, and as “$b$ is preferred to $a$ no more than $d$ is preferred to cc” when the converse is true (naturally, $(a,b)\dot{\succsim}_{l}(c,d)$ will also hold when $a$ is preferred to $b$ and $d$ is preferred to $c$). When no ambiguity arises, we will interpret $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ more succinctly as suggesting that the intensity difference between $a$ and $b$ is weakly larger than that between $c$ and $d$.
In line with these interpretations and the literature, agent $l$’s binary preference relation $\succsim_{l}$ on $A$ is derived from their intensity relation $\dot{\succsim}_{l}$ by $a\succsim_{l}b\Leftrightarrow(a,b)\,\dot{\succsim}_{l}\,(b,a)$. We will refer to $\succsim_{l}$ as the preference relation that is induced by $\dot{\succsim}_{l}$. As usual, the asymmetric and symmetric parts of $\succsim_{l}$ are denoted by $\succ_{l}$ and $\sim_{l}$. Under the structure that $\dot{\succsim}_{l}$ will be endowed with below, both this relation and its induced $\succsim_{l}$ are weak orders on their respective domains.
For $a,b\in A$, the intensity-equivalence class of $(a,b)\in A\times A$ under $\dot{\succsim}_{l}$ is defined by
$[a,b]$ $:=$ $\{(a^{\prime},b^{\prime})\in A\times A:(a^{\prime},b^{\prime})\,\dot{\sim}_{l}\,(a,b)\}$ (1)
Furthermore, the quotient set of $A\times A$ under $\dot{\sim}_{l}$ is defined by
$(A\times A)_{\dot{\sim}_{l}}$ $:=$ $\{[a,b]:[a,b]\text{ is an intensity-equivalence class under }\dot{\sim}_{l}\}$ (2)
The definitions that follow omit existential quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
Definition 1.
A relation $\dot{\succsim}_{l}$ has a utility-difference representation if there is a function $u_{l}:A\rightarrow\mathbb{R}$ that satisfies
$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $u_{l}(a)-u_{l}(b)\geq u_{l}(c)-u_{l}(d)$ (3)
In an influential article, psychologist S. S. Stevens (1946) distinguished between four scales of measurement that are afforded by some numerical assignment on a set of items: nominal (qualitative/categorical measurement); ordinal (invariance up to a strictly increasing transformation); interval (invariance up to positive affine transformation); and ratio (invariance up to a positive linear transformation). Utility-difference representations were added to this list by psychologist C. H. Coombs (1950) as those that correspond to ordered metric scale measurement. These are more general than neoclassical cardinal utility representations (Lange, 1934; Alt, 1936, 1971; Samuelson, 1938) because $u_{l}$ in the latter case is an interval scale whereas in the former it “falls logically between an interval scale and an ordinal scale” (p. 145). Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a subset of utility-difference comparisons whose order is indeed preserved by all ordinal transformations of the original representation. In our environment, whenever $u_{l}$ in (1) is not cardinal it has the complex uniqueness property of additive representations on finite sets.
Scott (1964), completing and extending the analysis that was initiated in Scott and Suppes (1958), characterized the existence of a utility-difference representation over a finite set by means of the following three axioms on $\dot{\succsim}_{l}$:
$(\mathbf{1_{\mathit{D}}})$ [Completeness]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ or $(c,d)\,\dot{\succsim}_{l}\,(a,b)$.
$(\mathbf{2_{\mathit{D}}})$ [Cancellation]. $(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
$(\mathbf{3_{\mathit{D}}})$ [Reversal]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ implies $(d,c)\,\dot{\succsim}_{l}\,(b,a)$.
The first imposes universal comparability on $\dot{\succsim}_{l}$. The third is essential for $\dot{\succsim}_{l}$ to be interpreted as a preference intensity relation rather than as an arbitrary preference relation over pairs. The second is known to have “no simple way of summarizing what it says about choice behavior” (Luce and Suppes, 1965, p. 277). Scott (1964, p. 244-5) explained how $(\mathbf{2_{\mathit{D}}})$ implies transitivity of $\dot{\succsim}_{l}$, but also remarked that $(\mathbf{2_{\mathit{D}}})$ is an infinite bundle of conditions (for each $t>0$, each $\pi,\sigma$), and it was shown in Scott and Suppes (1958) that no finite number of them could be sufficient”. Once the number $n$ of alternatives in $A$ is fixed, there are obviously finitely many such conditions to be satisfied. Scott’s remark, however, clarifies that it is impossible to reduce this condition to a finite number of statements/axioms that, together with $(\mathbf{1_{\mathit{D}}})$ and $(\mathbf{3_{\mathit{D}}})$, would be sufficient to characterize (3) for any finite $n$.
In Gerasimou (2021, 2022) this author recently proposed and characterized the class of ordinal intensity relations that admit the following more general numerical representation.
Definition 2.
A relation $\dot{\succsim}_{l}$ is representable by a preference intensity function if there is a mapping $s_{l}:A\times A\rightarrow\mathbb{R}$, unique up to a strictly increasing transformation, such that
$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(c,d)$ (4)
$(a,b)\,\dot{\succsim}_{l}\,(b,a)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(e,e)\geq s_{l}(b,a)$ (5)
$\min\{s_{l}(a,b),s_{l}(b,c)\}\geq s_{l}(e,e)$ $\Longrightarrow$ $s_{l}(a,c)\geq\max\{s_{l}(a,b),s_{l}(b,c)\},$ (6)
with a strict inequality on the left hand side of (6) implying a strict inequality on the right.
In words, $s_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)]. Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
$(\mathbf{4_{\mathit{D}}})$ [Translocation Consistency]. $(a,c)\,\dot{\succsim}_{l}\,(b,c)$ if and only if $(a,b)\,\dot{\succsim}_{l}\,(b,a)$.
We denote by $\mathcal{I}$ the class of intensity relations that admit this ordinal representation.
The (pseudo-)cardinal utility-difference model is nested by the ordinal model in the special case where “lateral consistency” [(6)] is strengthened to “additivity” [(7)]:
Lemma 1.
A relation $\dot{\succsim}_{l}$ that is represented by $s_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widehat{s}_{l}$ of $s_{l}$ satisfies
$\widehat{s}_{l}(a,c)$ $=$ $\widehat{s}_{l}(a,b)+\widehat{s}_{l}(b,c)$ (7)
Proof.
Upon viewing $\widehat{s}_{l}$ as a general function of two variables, (7) is known as Sincov’s functional equation, whose solution is some $f:A\rightarrow\mathbb{R}$, unique up to an additive constant, such that $\widehat{s}_{l}(a,b)\equiv f(a)-f(b)$ (Aczél, 1966). This establishes the “if” part. The “only if” part is immediate upon defining $\widehat{s}_{l}(a,b):=u_{l}(a)-u_{l}(b)$ for the postulated function $u_{l}$ with the requisite representation property. ∎
The normalization of $s_{l}$ introduced next is important for the rest of this paper.
Definition 3.
A preference intensity function $s_{l}$ represents $\dot{\succsim}_{l}$ canonically if
$s_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
with $k$ the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
Lemma 2.
Any relation $\dot{\succsim}_{l}\in\mathcal{I}$ admits a canonical preference intensity representation.
Proof.
Let $s_{l}$ be an arbitrary preference intensity function for $\dot{\succsim}_{l}\in\mathcal{I}$. Define the level set of $s_{l}$ at $(a,b)\in A\times A$ by $\mathbf{s}_{l}[a,b]:=\{(a^{\prime},b^{\prime})\in A\times A:s_{l}(a,b)=s_{l}(a^{\prime},b^{\prime})\}$. Further, define $r_{l}:A\times A\rightarrow\mathbb{R}$ by
$r_{l}(a,b)$ $:=$ $\left\{\begin{array}[]{ll}|\{\mathbf{s}_{l}[a^{\prime},b^{\prime}]:s_{l}(a,b)>s_{l}(a^{\prime},b^{\prime})\geq s_{l}(e,e)\}|,&\text{if }s_{l}(a,b)>s_{l}(e,e)\\ s_{l}(a,b),&\text{if }s_{l}(a,b)=s_{l}(e,e)\\ -r_{l}(b,a),&\text{if }s_{l}(a,b)<s_{l}(e,e)\end{array}\right$.
By construction, $r_{l}$ is canonical and $r_{l}(a,b)\geq r_{l}(c,d)\Leftrightarrow s_{l}(a,b)\geq s_{l}(c,d)$. ∎
We assume throughout that every $\dot{\succsim}_{l}$ belongs to $\mathcal{I}$, hence is representable canonically. Our subsequent analysis will further impose the following additional property on $\dot{\succsim}_{l}$:
$(\mathbf{5_{\mathit{D}}})$ [Strictness]. $(a,b)\,\mathord{\not\mathrel{\dot{\sim}_{l}}}\,(c,d)$ for all distinct off-diagonal pairs $(a,b),(c,d)\in A\times A$.
This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$. It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies.
As we elaborate in the next section, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$. In particular, the only non-trivial intensity-equivalence class that exists when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
Table 1: The number of distinct intensity relations of different classes for small values of $n$.
Strict & utility-difference Strict Strict ordinal
Strict representable ordinally representable intensity
$n=$ orders intensity relations intensity relations profiles
3 6 12 12 $12^{3}$
4 24 240 384 $384^{4}$
5 120 13,680 92,160 $92,160^{5}$
Finally, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit, respectively, a (pseudo-cardinal) utility-difference and an ordinal representation when $n\leq 5$, and was produced with Minizinc (Nethercote et al., 2007; Gecode Team, 2006).
3 The Comparability Postulate
[…]
Interpersonal Comparisons of Ordinal Intensities
Given an intensity profile $S\in\mathcal{\widehat{S}}$ and the corresponding canonical profile $s$, the statement
$s_{l}(a,b)$ $>$ $s_{m}(c,d)\;>\;0$ (10)
is interpreted as suggesting that agent $l$ prefers $a$ to $b$ more than agent $m$ prefers $c$ to $d$.
Towards defending this new concept we first recall that preference intensities at the level of the individual decision maker are not assumed here to be quantifiable beyond the level of an ordinal ranking. Yet some information about the generally different welfare effects of having $a$ or $b$ vs $c$ or $d$ is available here. Suppose, in particular, that agent $l$ prefers $a$ to $b$ to $c$ to $d$, and the first over the second more than the third over the fourth. This is equivalent to saying that her intensity difference at pair $(a,b)$ is greater than at $(c,d)$. Consider agent $m$ next, and suppose for simplicity that the same applies to him too: $s_{m}(a,b)>s_{m}(c,d)$. Considering this, and that both have preferences and intensities over the same set of alternatives, how might (10) be interpreted?
It is clearly a statement of fact that the intensity difference within pair $(a,b)$ is ranked higher by agent $l$ in her intensity ordering than the corresponding difference within pair $(c,d)$ is ranked by agent $m$ in his. Now recall that, by the Strictness assumption, $l$ and $m$ have the same number of possible rank positions in their respective intensity orderings which pairs $(a,b)$ and $(c,d)$ can occupy. In other words, the agents’ ordinal intensity scales coincide: if $(a,b)$ and $(c,d)$ are ranked 5th and 8th in $l$’s and $m$’s scales, respectively, then they are ranked 5th and 8th out of the same $k$ possible positions. As we mentioned in the Introduction, this logic resembles the way scoring rules such as the well-known Borda-count (Borda, 1781) normalization of ordinal utilities or scoring-like (see d’Aspremont, 2018) rules such as the relative-utilitarian (Dhillon and Mertens, 1999) normalization of cardinal utilities perform interpersonal comparisons in their respective environments.
The question now emerges: Should the different agents’ intensity orderings be treated equally by the social planner? While equal treatment may or may not be an appropriate approach to follow in practice, depending on what else is known about agents and the problem at hand, since the intensity orderings are assumed here to encode all available welfare-relevant information, treating them in any way other than equal would call for a justification that appears elusive. Against this background, we view the suggested interpretation of (10) as a reasonable starting point for interpersonal comparisons in this informational environment.
On the analytical side, another relevant question also arises: in the special case where the ordinal intensity ordering of every agent is defined by how the pairwise differences of a cardinal utility function are ordered, and these differences in turn define a strict intensity profile, does interpersonal comparisons of ordinal intensities reduce to the interpersonal comparability of utility differences that is required by utilitarian aggregation? As the next example shows, the answer is no.
Example 1.
Let $n=3$ and consider the utility profile $U:=\big{(}(53,38,9),(46,33,21),$ $(20,30,50)\bigr{)}$ over $A=\{a_{1},a_{2},a_{3}\}$, where each $u_{l}$ in $U$, $l\leq 3$, is constructed so its values add up to 100 (this is without loss). It is readily seen that the intensity profile defined by the intra-personal utility differences in $U$ is in $\mathcal{\widehat{S}}$. Also, defining $\dot{\succ}_{i}$ by $(a_{i},a_{j})\dot{\succ}_{l}(a_{k},a_{m})$ $\Leftrightarrow$ $u_{l}(a_{i})-u_{l}(a_{j})>u_{l}(a_{k})-u_{l}(a_{m})$, one observes that $(a_{2},a_{3})\,\dot{\succ}_{1}\,(a_{1},a_{2})$ and $(a_{1},a_{2})\,\dot{\succ}_{2}\,(a_{2},a_{3})$ hold, hence $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$ too. However, $u_{1}(a_{1})-u_{1}(a_{2})>u_{2}(a_{1})-u_{2}(a_{2})$ is also true. $\blacklozenge$
Thus, even if there is some ordinal transformation $\widehat{s}_{i}$ of the canonical intensity function $s_{i}$ such that $\widehat{s}_{i}(a,b)\equiv u_{i}(a)-u_{i}(b)$ for some cardinally unique $u_{i}$ [cf. (7)], the above clarifies that (10) neither implies nor is implied by the inequality $u_{l}(a)-u_{l}(b)>u_{m}(c)-u_{m}(d)$. For this reason, and in light of the preceding discussion, in such rich special environments we must generally distinguish between the relative intensity difference that is captured by $s_{l}(a,b)>s_{m}(c,d)$ and the absolute intensity difference that is captured by the above interpersonal utility-difference inequality, which is preserved under any positive affine transformation of agents’ utilities that involves a common multiplicative factor (i.e. the Cardinal Unit Comparability informational basis—see also Section 4.2).
4 Intensity Dominance and Intensity-Efficient Allocations
4.1 Definitions
Since the number of agents and items are assumed to coincide, an allocation of the $n$ goods is a permutation on $A$. The set of all allocations is denoted by $\mathcal{A}$. As we discussed in the Introduction, we are interested in the assignment of the $n$ objects in $A$ to the $n$ agents in $V$ in a way that satisfies Pareto-efficiency but also improves upon it in delivering a narrower set of acceptable allocations that satisfy some additional desideratum that is appealing from a distributive-justice perspective and is operational in the present analytical environment where information on ordinal preference intensities and interpersonal comparisons thereof is assumed to be available. To this end, we proceed with the introduction of the following novel notions of dominance and efficiency.
Definition 4.
Let $S=\big{(}\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n}\bigr{)}\in\widehat{\mathcal{S}}$ be a strict intensity profile and $s$ its canonical representation. Given allocations $x$ and $y$, the former is said to intensity-dominate the latter if, for every pair of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$ it holds that $s_{i}(x_{i},x_{j})\geq s_{j}(y_{j},y_{i})$, and there is at least one pair $(i,j)$ where this inequality is strict.
Definition 5.
An allocation is intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ if it is Pareto-efficient at the induced preference profile $P_{S}\in\mathcal{P}^{V}$ and is not intensity-dominated by another Pareto-efficient allocation at $P_{S}$. It is intensity-dominant if it intensity-dominates every other Pareto-efficient allocation at $P_{S}$.
Definition 6.
Two allocations $x$ and $y$ that are intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ are equivalent if $s_{i}(x_{i},x_{j})=s_{j}(y_{j},y_{i})$ for all pairs of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$, and $x_{k}=y_{k}$ for every other agent $k\leq n$.
If $x$ and $y$ are Pareto-efficient allocations and $x$ intensity-dominates $y$, then in every pair of agents that is “flipped” by $x$ and $y$ in the sense that both allocations assign the same two alternatives $a$ and $b$ to the two agents in that pair but do so in opposite ways, the agent receiving $a$ (which, under the postulated Pareto-efficiency, is the mutually preferred alternative) under $x$ prefers it to $b$ weakly more than the agent receiving it under $y$ and, in at least one case, strictly more. An allocation that is both Pareto efficient and not intensity-dominated is intensity-efficient. Two intensity-efficient allocations $x$ and $y$ are equivalent if the agents in every “flipping” pair $(i,j)$ who are assigned $a$ and $b$ by $x$ and $y$ but in opposite ways have the same preference intensity for $a$ over $b$ in the sense that the intensity difference at $(a,b)$ is ranked the same way in the two agents’ intensity orderings, while every agent who is not in such a pair is treated identically by $x$ and $y$.
4.2 Intensity-Efficient and Utilitarian Allocations Are Distinct
To enable a comparison between the concept introduced above and the benchmark Pareto-efficiency refinement afforded by classical utilitarianism, we assume throughout this subsection that, in addition to satisfying the strictness condition $(\mathbf{5_{\mathit{D}}})$, each $\dot{\succsim}_{l}$ is utility-difference representable as in (3) under some cardinal utility index $u_{l}$. This is the case, for example, when $u_{l}:A\rightarrow\mathbb{R}$ is the utility function associated with an expected-utility representation of preferences over lotteries over $A$ and the intensity order defined by the $u_{l}$-differences is strictly convex or strictly concave relative to the agent’s induced (strict) preferences. With this assumption in place, we finally invoke the Cardinal Unit Comparability (Roemer, 1996; Baccelli, 2023) informational basis to social welfare. More specifically, we assume that: (i) $u_{l}:A\rightarrow\mathbb{R}$ is unique up to a positive affine transformation and features utility differences that define an intensity relation $\dot{\succsim}_{l}$ that abides with $(\mathbf{5_{\mathit{D}}})$; (ii) interpersonal utility-difference comparisons are possible under such transformations of any utility profile $U=(u_{1},\ldots,u_{n})$ where each $u_{l}$ is of this kind, as long as the scale/multiplicative parameter of such transformations is the same for all $u_{l}$. As is well-known, this informational basis is sufficient [^1] The reason why it is not necessary is closely related to the complex uniqueness properties of general utility-difference representations in the sense of (3). See Bossert (1991) for more on this. for the classical utilitarian social welfare functional—recalled below—to be meaningful.
Fixing a utility profile $U$ with the above properties, we refer to it as strict if the ordering of every agent’s utility differences is strict (see Example 1). In this case, clearly, the ordinal intensity profile $S$ that is induced by $U$ is also strict, i.e. an element of $\mathcal{\widehat{S}}$. Finally, recall that $x_{cu}(U)\in\mathcal{A}$ is a classical utilitarian allocation under profile $U$ if it holds that
$x_{cu}(U)$ $\in$ $\arg\max_{y\in\mathcal{A}}\sum_{l=1}^{n}u_{l}(y_{l})$ (11)
Proposition 1.
The sets of intensity-efficient and classical utilitarian allocations associated with strict utility profiles are generally distinct.
Proof.
Consider the strict utility profile of Example 1. Observe that the Pareto-efficient allocations are $(a_{1},a_{2},a_{3})$ and $(a_{2},a_{1},a_{3})$, and that the former allocation is the utilitarian one. Yet this is intensity-dominated by the latter since $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$. ∎
This potential divergence between intensity efficiency and the hallmark cardinal-welfarist allocation criterion is important because it shows that the hereby proposed notion is indeed distinct even under the richest possible informational assumptions. Furthermore, while perhaps not obvious ex ante, as is clarified by Example 1 and the discussion at the end of in Section 3, the reason why such a divergence may occur is that the classical utilitarian criterion is guided by the utility differences across agents. By contrast, in this special environment where intensity orderings are defined by the relative ordering rather than the absolute magnitude of utility differences, the Interpersonal Comparisons of Ordinal Intensities postulate upon which the intensity efficiency criterion is built compares the relative ranking of those differences across agents’ intensity orderings.
Naturally, one may inquire about the conditions under which classical utilitarian allocations are intensity-efficient. To this end, and motivated by the preceding remarks, we introduce the following condition that a utility profile may satisfy, which is sufficient—though not necessary—for utilitarian allocations to be intensity-efficient.
Definition 7.
A utility profile $U=(u_{1},\ldots,u_{n})$ is balanced if, for any two agents $i,j$ and alternatives $a,b$,
$u_{i}(a)-u_{i}(b)>u_{j}(a)-u_{j}(b)$ $\Longleftrightarrow$ $s_{i}(a,b)>s_{j}(a,b),$ (12)
where $s=(s_{1},\ldots,s_{n})$ is the profile of canonical preference intensity functions that represent the profile $S=(\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n})\in\widehat{\mathcal{S}}$ that is defined by the utility differences in $U$.
In words, a cardinal utility profile is balanced if interpersonal differences in the levels of utility differences are perfectly aligned with interpersonal differences in their rank orders.
Proposition 2.
Classical utilitarian allocations of strict and balanced utility profiles are intensity-efficient.
Proof.
Under the postulated conditions, suppose to the contrary that $x(U)$ is a utilitarian but not intensity-efficient allocation. Then, given that agent $l$’s intensity relation $\dot{\succsim}_{l}$ is defined by $(a,b)\,\dot{\succsim}_{l}\,(c,d)\Leftrightarrow u_{l}(a)-u_{l}(b)\geq u_{l}(c)-u_{l}(d)$, from the definition of intensity-dominance and intensity-efficiency there exist at least one pair of agents $(i,j)$ and of objects $(a,b)$ such that $x_{i}(U)=a$, $x_{j}(U)=b$ and $s_{j}(a,b)>s_{i}(a,b)$. By balancedness, this implies
$u_{j}(a)-u_{j}(b)$ $>$ $u_{i}(a)-u_{i}(b)$ (13)
Consider allocation $x^{\prime}$ that is identical to $x(U)$ except that $x^{\prime}_{i}=b$ and $x^{\prime}_{j}=a$. From (13) we get $u_{j}(a)-u_{i}(a)>u_{j}(b)-u_{i}(b)$. This shows that the utilitarian benefit from the $i\leftrightarrow j$ swap in the transition from $x(U)$ to $x^{\prime}$, $u_{j}(a)-u_{i}(a)$, exceeds the corresponding loss, $u_{j}(b)-u_{i}(b)$. It follows that $\sum_{l=1}^{n}u_{l}(x^{\prime}_{l})>\sum_{l=1}^{n}u_{l}\big{(}x_{l}(U)\bigr{)}$, which contradicts the postulated utilitarian-efficiency of $x(U)$. ∎
It is well-known that utilitarian allocations are Pareto-efficient. Proposition 2 shows that, when the utility profile is strict and balanced, this benchmark criterion has the additional normative status conveyed by intensity-efficiency. Furthermore, Proposition 2 implies the following existence result that is perhaps worth stating explicitly:
Corollary 1.
An intensity-efficient allocation exists for every intensity profile that is induced by a strict and balanced utility profile.
4.3 Existence without Cardinality Assumptions
We now return to the general environment in which intensity profiles come from the class of relations that were specified in Section 4.1, without imposing any further restrictions such as cardinal or even utility-difference representability of the individuals’ intensities.
Theorem 1.
An intensity-efficient allocation exists for every profile $S\in\mathcal{\widehat{S}}$ when $n=3$ but not in general when $n\geq 4$.
Proof.
The argument proceeds by considering the possible ways in which an arbitrary profile $S\in\mathcal{\widehat{S}}$ when $n=3$ might generate a sequence of distinct Pareto-efficient allocations that are implicated in an intensity-dominance cycle. To this end, let $D$ be the intensity-dominance relation that is introduced in Definition 4. Suppose to the contrary that
$w^{1}Dw^{2}D\ldots Dw^{k}Dw^{1}$ (14)
for Pareto efficient allocations $w^{1},\ldots,w^{k}$ on $X:=\{a,b,c\}$.
Observation 1.
$n=3$ implies that for any two allocations $w^{i},w^{i+1}$ such that $w^{i}Dw^{i+1}$ it must be that $w^{i}_{l}=w^{i+1}_{l}$ for exactly one agent $l\in\{1,2,3\}$ and $(w^{i}_{j},w^{i}_{k})=(w^{i+1}_{k},w^{i+1}_{j})$ for $j,k\neq l$.
Observation 2.
$n=3$ implies $k\leq 6$.
Observation 3.
Pareto efficiency of $w^{i}=(a^{\prime},b^{\prime},c^{\prime})$ and Strictness together imply
$s_{2}(a^{\prime},b^{\prime})>0$ $\Longrightarrow$ $s_{1}(a^{\prime},b^{\prime})>0$ (15)
$s_{3}(b^{\prime},c^{\prime})>0$ $\Longrightarrow$ $s_{2}(b^{\prime},c^{\prime})>0$ (16)
$s_{3}(a^{\prime},c^{\prime})>0$ $\Longrightarrow$ $s_{1}(a^{\prime},c^{\prime})>0$ (17)
Observation 4.
Strictness and canonicality of $s_{i}$ imply $s_{i}(a^{\prime},b^{\prime})=s_{i}(c^{\prime},d^{\prime})>0\Leftrightarrow(a^{\prime},b^{\prime})=(c^{\prime},d^{\prime})$ and, jointly with $n=3$, also $s_{i}(a^{\prime},b^{\prime})>0\Leftrightarrow s_{i}(a^{\prime},b^{\prime})\in\{1,2,3\}$.
Notice that (14) is impossible for $k=2$ because $D$ is asymmetric by construction. Suppose $k=3$. Without loss of generality, write $w^{1}:=(a,b,c)$ and $w^{2}:=(b,a,c)$. Then, by Observation 1, either $w^{3}=(b,c,a)$ or $w^{3}=(c,a,b)$. Since, in both cases, $w^{1}$ and $w^{3}$ are $D$-incomparable by construction, the $w^{3}Dw^{1}$ postulate in (14) is contradicted.
Now suppose $k=4$. By (14) and the above implications, we may take $w^{1}$, $w^{2}$, $w^{3}$ to be as in the $k=3$ case, from which it then follows that allocation $w^{4}$ must be either $(c,b,a)$ or $(a,c,b)$. Notice that either possibility is compatible with $w^{3}=(b,c,a)$ and with $w^{3}=(c,a,b)$. We therefore have the following 4 cases to consider:
[[Note: in what follows we make repeated use—often without explicit reference—of the Pareto-efficiency implications (15)–(17), the skew-symmetry property $s_{l}(a,b)=-s_{l}(b,a)$ of canonical representations and, whenever exact values of the $s_{i}$ functions are asserted, of the lateral-consistency property (6) together with the assumption that every $s_{i}$ is canonical and strict (cf Observations 34).]]
Case 1. $w^{3}=(b,c,a)$, $w^{4}=(c,b,a)$. By the definition of $D$, and by the above assumptions:
$w^{1}Dw^{2}$ $\Longrightarrow$ $s_{1}(a,b)>s_{2}(a,b)$ (18)
$w^{2}Dw^{3}$ $\Longrightarrow$ $s_{2}(a,c)>s_{3}(a,c)$ (19)
$w^{3}Dw^{4}$ $\Longrightarrow$ $s_{1}(b,c)>s_{2}(b,c)$
$w^{4}Dw^{1}$ $\Longrightarrow$ $s_{3}(a,c)>s_{1}(a,c)$
Therefore,
$s_{2}(a,c)>s_{3}(a,c)>s_{1}(a,c)$ (20)
Since $w^{i}$ is Pareto efficient for $i\leq 4$, it follows from (15)–(17) that there are 4 subcases to consider:
Subcase 1-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{2}(b,c)>0$. By (6) and the fact that $s_{1},s_{2}$ are canonical, this implies $s_{1}(a,c)=3$, which contradicts (20).
Subcase 1-ii. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{2}(c,b)>s_{1}(c,b)>0$. Suppose $s_{1}(a,c)>0$ is also true. Then, by (6), (20) and the fact that $s_{1}$ is canonical, $s_{1}(a,c)=1$ and $s_{1}(a,b)=3$. If $s_{2}(a,c)>0$ is also true, then $s_{2}(a,b)=3$. This contradicts $s_{1}(a,b)>s_{2}(a,b)$. So, it must be that $s_{2}(c,a)>0$ instead. But in this case $s_{2}(c,a)>0$, $s_{2}(a,b)>0$ and (6) together imply $s_{2}(c,b)=3$. This contradicts $s_{2}(a,c)=3$ which is now implied by (20) and the fact that the profile $s$ is canonical. Thus, it must be that $s_{1}(c,a)>0$ instead. So now we have $s_{1}(c,a)>0$, $s_{1}(a,b)>0$, which implies $s_{1}(c,b)=3$. But since, by assumption, $s_{2}(c,b)>s_{1}(c,b)$ and $s_{2}$ is canonical, this is a contradiction.
Subcase 1-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{2}(b,c)>0$. Suppose first that $s_{2}(a,c)>0$ is also true. Then, $s_{2}(b,a)>0$ and $s_{2}(a,c)>0$ implies $s_{2}(b,c)=3$. If $s_{1}(a,c)>0$ is also true, then (20) and the fact that $s$ is canonical together imply $s_{2}(a,c)=s_{2}(b,c)$, which contradicts (6) and Strictness. So, it must be that $s_{1}(c,a)>0$. From $s_{1}(b,c)>0$ and $s_{1}(c,a)>0$ we now get $s_{1}(b,a)=3$. In view of $s$ being canonical, this contradicts $s_{2}(b,a)>s_{1}(b,a)$.
Subcase 1-iv. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{2}(c,b)>s_{1}(c,b)>0$. Because $s$ is canonical, this and (6) readily imply $s_{1}(c,a)=s_{2}(c,a)=3$. But since (20) is equivalent to $s_{1}(c,a)>s_{3}(c,a)>s_{2}(c,a)$, this is a contradiction.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$.
Case 2. $w^{3}=(c,a,b)$, $w^{4}=(c,b,a)$.
We again have $w^{1}Dw^{2}\Longrightarrow s_{1}(a,b)>s_{2}(a,b)$ and $w^{2}Dw^{3}\Longrightarrow s_{1}(b,c)>s_{3}(b,c)$ and, in addition,
$w^{3}Dw^{4}$ $\Longrightarrow$ $s_{2}(a,b)>s_{3}(a,b)$ (21)
$w^{4}Dw^{1}$ $\Longrightarrow$ $s_{3}(a,c)>s_{1}(a,c)$ (22)
Therefore,
$s_{1}(a,b)>s_{2}(a,b)>s_{3}(a,b)$ (23)
In view of (15)–(17), we can now consider the following 4 possible subcases:
Subcase 2-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. By (23) and (15)–(17), the former postulate implies $s_{3}(a,b)>0$. Since $s$ is canonical, this further implies $s_{3}(a,b)=1$, $s_{2}(a,b)=2$ and $s_{1}(a,b)=3$. This, canonicality of $s$ and $s_{1}(b,c)>s_{3}(b,c)>0$ together imply $s_{1}(b,c)=2$ and $s_{3}(b,c)=1=s_{3}(a,b)$, which contradicts Strictness.
[…]
Subcase 2-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The former postulate, together with (23) and canonicality, implies $s_{3}(b,a)=3$, $s_{2}(b,a)=2$, $s_{1}(b,a)=1$. By (21), either $s_{3}(a,c)>s_{1}(a,c)>0$ or $s_{1}(c,a)>s_{3}(c,a)>0$ also holds. Consider the first possibility. From $s_{1}(b,a)=1$, $s_{1}(a,c)>0$, (6) and canonicality we get $s_{1}(a,c)=2$. This and (22) implies $s_{3}(a,c)=3$. Since $s_{3}(b,a)=3$ is also true, this contradicts Strictness. Hence, it must be that $s_{1}(c,a)>s_{3}(c,a)>0$. But in this case $s_{1}(b,c)>0$, $s_{1}(c,a)>0$, (6) and canonicality imply $s_{1}(b,a)=3$, which contradicts (23).
Subcase 2-iv. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{3}(c,b)>s_{1}(c,b)>0$. As in 2-iii, we have $s_{3}(b,a)=3$, $s_{2}(b,a)=2$, $s_{1}(b,a)=1$. But $s_{3}(c,b)>0$, $s_{3}(b,a)>0$ and (6) imply $s_{3}(c,a)>s_{3}(b,a)=3$ which, by canonicality, is impossible.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$ too.
Case 3. $w^{3}=(b,c,a)$, $w^{4}=(a,c,b)$.
We now have
$w^{3}Dw^{4}$ $\Longrightarrow$ $s_{3}(a,b)>s_{1}(a,b)$ (24)
$w^{4}Dw^{1}$ $\Longrightarrow$ $s_{3}(b,c)>s_{2}(b,c)$ (25)
It follows that
$s_{3}(a,b)>s_{1}(a,b)>s_{2}(a,b)$ (26)
We consider the 4 subcases that are now possible:
Subcase 3-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The latter, together with (15)–(17), (25) and canonicality, implies $s_{3}(a,b)=1$. But canonicality, (15)–(17) and (17) also implies $s_{3}(a,b)=3$, which contradicts Strictness.
Subcase 3-ii. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{3}(c,b)>s_{1}(c,b)>0$. The first postulate and (26), together with canonicality, implies $s_{3}(a,b)=3$, $s_{1}(a,b)=2$ and $s_{2}(a,b)=1$. Since $s_{3}(c,b)>s_{1}(c,b)>0$ is also assumed, this and canonicality further imply $s_{3}(c,b)=2$. Now, because $s_{3}(a,b)>0$ and $s_{3}(c,b)>0$, it follows from (6) that $s_{3}(c,a)>0$ too. But (6) in this case further implies $s_{3}(c,a)>s_{3}(a,b)=3$, which is impossible.
Subcase 3-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The first postulate, together with (26) and canonicality, implies $s_{2}(b,a)=3$, $s_{1}(b,a)=2$, $s_{3}(b,a)=1$. The second postulate and $s_{1}(b,a)=2$, together with Strictness, implies $s_{1}(b,c)=3$. This in turn implies $s_{1}(a,c)=1$ or $s_{1}(c,a)=1$. If the latter is true, then $s_{1}(b,c)>0$, $s_{1}(c,a)>0$ and (6), together with canonicality, implies $s_{1}(b,a)=3$, a contradiction. Hence, it must be that $s_{1}(a,c)=1$. We therefore have $s_{1}(b,a)=2$, $s_{1}(a,c)=1$ and, by (6) and canonicality, $s_{1}(b,c)=3$. From (15)–(17), (19), (25) and canonicality we also know that $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)>0$ implies $s_{3}(b,c)=2$ and $s_{2}(b,c)=1$. Thus, we have $s_{3}(b,a)>0$, $s_{3}(b,c)>0$ and, by (15)–(17) and $s_{1}(a,c)>0$, also $s_{3}(a,c)>0$. But $s_{3}(b,a)>0$, $s_{3}(a,c)>0$ together with (6) and canonicality implies $s_{3}(b,c)=3$, a contradiction.
Subcase 3-iv. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{3}(c,b)>s_{1}(c,b)>0$. These readily imply $s_{1}(c,a)=3$. As above, (26) implies $s_{2}(b,a)=3$, $s_{1}(b,a)=2$ and $s_{3}(b,a)=1$. By (6), $s_{1}(c,a)=3$ and $s_{1}(a,b)=2$ implies $s_{1}(c,b)=1$. From the above postulates and from (25), $s_{2}(c,b)>s_{3}(c,b)>s_{1}(c,b)$ further implies $s_{2}(c,b)=3$, which contradicts $s_{2}(b,a)=3$ and Strictness.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$ here as well.
Case 4. $w^{3}=(c,a,b)$, $w^{4}=(a,c,b)$. It is now true that
$w^{3}Dw^{4}$ $\Longrightarrow s_{2}(a,c)>s_{1}(a,c)$ (27)
$w^{4}Dw^{1}$ $\Longrightarrow s_{3}(b,c)>s_{2}(b,c)$ (28)
which, together with (18)–(19), imply
$s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)$ (29)
Suppose first that $s_{2}(b,c)>0$. Then, by (29) and (15)–(17), $s_{1}(b,c)=3$, $s_{2}(b,c)=1$ and $s_{3}(b,c)=2$. From $s_{1}(b,c)=3$ and (6) we also get $s_{1}(b,a)>0$ and $s_{1}(a,c)>0$. Hence, by (18), $s_{2}(b,a)>s_{1}(b,a)>0$ and, by (27), $s_{2}(a,c)>s_{1}(a,c)>0$. These inequalities and (6) together imply $s_{2}(b,c)=3$, which is a contradiction.
Now suppose instead that $s_{2}(b,c)<0$, i.e. $s_{2}(c,b)>0$. It follows from (29) that $s_{2}(c,b)=3$, $s_{3}(c,b)=2$ and $s_{1}(c,b)=1$. Suppose $s_{1}(a,c)>0$. From (27) and (15)–(17), $s_{2}(a,c)>0$. Since $s_{2}(a,c)>0$ and $s_{3}(c,b)>0$, by (6) we get $s_{2}(a,b)>s_{2}(c,b)=3$, which is impossible. Hence, $s_{1}(c,a)>0$ holds instead and, from (27) and (15)–(17), $s_{1}(c,a)>s_{2}(c,a)>0$ is also true. Suppose $s_{2}(a,b)>0$ holds too. By (18), $s_{1}(a,b)>0$. By (6) and $s_{1}(c,a)>0$, $s_{1}(a,b)>0$ we get $s_{1}(c,b)=3$, a contradiction. Hence, $s_{2}(b,a)>0$ must be true instead and, by (18), $s_{2}(b,a)>s_{1}(b,a)>0$ also. So, we have $s_{2}(c,b)>0$ and $s_{2}(b,a)>0$, which, by (6), implies $s_{2}(c,a)>s_{(}c,b)=3$. This too is a contradiction.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$ also.
Next, suppose $k=5$. Arguing as above, allocations $w^{1},\ldots,w^{4}$ in (14) must be as in one of the four cases considered previously. Combined with the fact that each $w^{i}$ in sequence $(w^{1},\ldots,w^{5})$ must be distinct and the notational convention $w^{1}=(a,b,c)$ and $w^{2}=(b,a,c)$, this gives rise to the following four possibilities:
[…]
Clearly, because either $w^{5}=(b,c,a)$ or $w^{5}=(c,a,b)$ must hold in all four cases, and recalling that $w^{1}=(a,b,c)$, it cannot be that $w^{5}Dw^{1}$.
Finally, suppose $k=6$. With allocations $w^{1},\ldots,w^{5}$ in (14) being as in the $k=5$ case that was just considered above, $w^{6}$ can only coincide with allocation $(a,c,b)$ in each of the four relevant cases. In view of the previous steps, these are as follows:
Case 1: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(b,c,a)$, $w^{4}=(c,b,a)$, $w^{5}=(c,a,b)$, $w^{6}=(a,c,b)$. By definition of $D$, and by the above assumptions: $w^{1}Dw^{2}\Longrightarrow s_{1}(a,b)>s_{2}(a,b)$; $w^{2}Dw^{3}\Longrightarrow s_{2}(a,c)>s_{3}(a,c)$; $w^{3}Dw^{4}\Longrightarrow s_{1}(b,c)>s_{2}(b,c)$; $w^{4}Dw^{5}\Longrightarrow s_{3}(a,b)>s_{2}(a,b)$; $w^{5}Dw^{6}\Longrightarrow s_{2}(a,c)>s_{1}(a,c)$; and $w^{6}Dw^{1}\Longrightarrow s_{3}(b,c)>s_{2}(b,c)$. It follows that
$s_{2}(a,c)>s_{1}(a,c)>s_{1}(a,b)>s_{2}(a,b)$ (30)
Suppose $s_{2}(a,b)>0$. Then, (30) implies $s_{2}(a,c)>3$, which contradicts canonicality. If $s_{2}(b,a)>0$ instead, then (30) together with skew-symmetry of $s_{2}$ implies $s_{2}(b,a)>3$ and results in the same contradiction.
Case 2: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(c,a,b)$, $w^{4}=(c,b,a)$, $w^{5}=(b,c,a)$, $w^{6}=(a,c,b)$. Notice that the postulated dominance implications $w^{1}Dw^{2}\Longrightarrow s_{1}(a,b)>s_{2}(a,b)$, $w^{3}Dw^{4}\Longrightarrow s_{2}(a,b)>s_{3}(a,b)$ and $w^{5}Dw^{6}\Longrightarrow s_{3}(a,b)>s_{1}(a,b)$ lead to $s_{1}(a,b)>s_{2}(a,b)>s_{3}(a,b)>s_{1}(a,b)$, which is absurd.
Case 3: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(b,c,a)$, $w^{4}=(a,c,b)$, $w^{5}=(c,a,b)$, $w^{6}=(a,c,b)$. Observe here that the postulated dominance implications $w^{4}Dw^{5}\Longrightarrow s_{1}(a,c)>s_{2}(a,c)$ and $w^{5}Dw^{6}\Longrightarrow s_{2}(a,c)>s_{1}(a,c)$ directly contradict each other.
Case 4: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(c,a,b)$, $w^{4}=(a,c,b)$, $w^{5}=(b,c,a)$, $w^{6}=(a,c,b)$. As in Case 2, the postulated dominance implications $w^{4}Dw^{5}\Longrightarrow s_{1}(a,b)>s_{3}(a,b)$ and $w^{5}Dw^{6}\Longrightarrow s_{3}(a,b)>s_{1}(a,b)$ result in the same contradiction.
It has therefore been shown that $D$ is acyclic when $n=3$, establishing the existence of an intensity-efficient allocation in this case.
For the claimed potential non-existence when $n\geq 4$, consider the example strict intensity profile on $A=\{a,b,c,d\}$ whose canonical representation is shown in the table below:
$s_{i}(a,b)=$ $i=$ 1 2 3 4
6 $(a,d)$ $(a,d)$ $(a,d)$ $(a,d)$
5 $(b,d)$ $(a,c)$ $(a,c)$ $(b,d)$
4 $(a,c)$ $(b,d)$ $(b,d)$ $(a,c)$
3 $(b,c)$ $(b,c)$ $(b,c)$ $(b,c)$
2 $(a,b)$ $(c,d)$ $(c,d)$ $(a,b)$
1 $(c,d)$ $(a,b)$ $(a,b)$ $(c,d)$
We start by observing that $s_{1}=s_{4}$, $s_{2}=s_{3}$ and $a\succ_{i}b\succ_{i}c\succ_{i}d$, $i=1,\ldots,4$. This implies that all 24 possible allocations, defined and listed below, are Pareto efficient.
$\begin{array}[]{llllllllllllllllll}x^{1}&=&(a,b,c,d)&&&x^{2}&=&(a,b,d,c)&&&x^{3}&=&(a,c,b,d)&&&x^{4}&=&(a,c,d,b)\\ x^{5}&=&(a,d,b,c)&&&x^{6}&=&(a,d,c,b)&&&x^{7}&=&(b,a,c,d)&&&x^{8}&=&(b,a,d,c)\\ x^{9}&=&(b,c,a,d)&&&x^{10}&=&(b,c,d,a)&&&x^{11}&=&(b,d,a,c)&&&x^{12}&=&(b,d,c,a)\\ x^{13}&=&(c,a,b,d)&&&x^{14}&=&(c,a,d,b)&&&x^{15}&=&(c,b,a,d)&&&x^{16}&=&(c,b,d,a)\\ x^{17}&=&(c,d,a,b)&&&x^{18}&=&(c,d,b,a)&&&x^{19}&=&(d,a,b,c)&&&x^{20}&=&(d,a,c,b)\\ x^{21}&=&(d,b,a,c)&&&x^{22}&=&(d,b,c,a)&&&x^{23}&=&(d,c,a,b)&&&x^{24}&=&(d,c,b,a)\end{array}$
The following comparisons, whose validity can be readily established by the reader, demonstrate that for each $x^{i}$, $i=1,\ldots,24$, in this set there is a distinct $x^{j}$ that intensity-dominates $x^{i}$—a situation denoted by $x^{j}Dx^{i}$.
$\begin{array}[]{llllllll}x^{6}Dx^{1}&x^{1}Dx^{2}&x^{4}Dx^{3}&x^{13}Dx^{4}&x^{3}Dx^{5}&x^{17}Dx^{6}&x^{1}Dx^{7}&x^{2}Dx^{8}\\ x^{3}Dx^{9}&x^{8}Dx^{10}&x^{9}Dx^{11}&x^{11}Dx^{12}&x^{18}Dx^{13}&x^{13}Dx^{14}&x^{17}Dx^{15}&x^{22}Dx^{16}\\ x^{23}Dx^{17}&x^{24}Dx^{18}&x^{8}Dx^{19}&x^{22}Dx^{20}&x^{11}Dx^{23}&x^{12}Dx^{22}&x^{24}Dx^{23}&x^{10}Dx^{24}\end{array}$
Therefore, no intensity-efficient allocation exists in this profile. ∎
The somewhat surprising fact that a plausible dominance concept may by cyclic and prevent an optimal entity to emerge invites an informal analogy between intensity-dominance cycles over allocations with at least 4 agents and Condorcet cycles over alternatives in pairwise majority-based preference aggregation with at least 3 agents (Condorcet, 1785). Importantly, however, although intensity-dominance cycles here may prevent refining the Pareto set, unlike the Condorcet social-welfare theoretic framework these do not lead to a “policy paralysis” problem because the Pareto set is always non-empty and, absent any distributively juster suggestions, one of them might be promoted by the social planner. That said, it is naturally of interest to understand better the conditions under which our proposed refinement of Pareto-efficiency is well-defined. This is the focus of the next subsection.
4.4 Existence Condition for General $n$: Homeo-Monotonicity
For any set $A^{\prime}\subseteq A$ and preference relation $\succ_{i}$ that is induced by an intensity relation $\dot{\succsim}_{i}\in\mathcal{I}$ we write $\succ_{i}^{A^{\prime}}$ for the restriction of $\succ_{i}$ on $A^{\prime}$.
Definition 8.
Given a profile $S\in\mathcal{\widehat{S}}$, a set of agents $V^{\prime}\subseteq V$ is said to be:
  1. 1.
    Envy-free with respect to a set of items $A^{\prime}\subseteq A$ if each $v_{i}\in V^{\prime}$ has a distinct $\succ_{i}$-best item in $A$, and this item (denoted $b_{i}^{\prime}$) belongs to $A^{\prime}$.
  2. 2.
    Top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ if $\succ_{i}^{A^{\prime}}=\succ_{j}^{A^{\prime}}$ for all $v_{i},v_{j}\in V^{\prime}$ and, when $A^{\prime}\subset A$, $a^{\prime}\in A^{\prime}$, $a\in A\setminus A^{\prime}$ holds, then $a^{\prime}\succ_{i}a$ for all $v_{i}\in V^{\prime}$.
An envy-free set $V^{\prime}$ with respect to $A^{\prime}$ consists of agents in $V^{\prime}$ who have distinct overall-best items in $A^{\prime}$. This terminology is motivated by the fact that, if $V^{\prime}\equiv V$ and $A^{\prime}\equiv A$, then each agent receiving their most preferred option results in an allocation that is Pareto-efficient and also envy-free (Varian, 1974; Moulin, 2004). A top-preference-coincident pair consists of agents whose preferences over $A^{\prime}$ are identical and also such that every item in $A^{\prime}$ is preferred to everything in $A\setminus A^{\prime}$, with no restrictions on preferences in the latter set.
Definition 9.
A set of agents $V^{\prime}\subseteq V$ that is top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ at profile $S\in\mathcal{\widehat{S}}$ is monotonic with respect to $A^{\prime}$ if $|V^{\prime}|=|A^{\prime}|$ and
$s_{1}(a_{1},a_{2})$ >> $s_{1}(a_{2},a_{3})$ >> $\cdots$ >> $s_{1}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
$\geq$ $\geq$ $\cdots$ $\geq$
$s_{2}(a_{1},a_{2})$ >> $s_{2}(a_{2},a_{3})$ >> $\cdots$ >> $s_{2}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
$\geq$ $\geq$ $\cdots$ $\geq$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
$\geq$ $\geq$ $\cdots$ $\geq$
$s_{|V^{\prime}|}(a_{1},a_{2})$ >> $s_{|V^{\prime}|}(a_{2},a_{3})$ >> $\cdots$ >> $s_{|V^{\prime}|}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
It is strictly monotonic if every weak inequality above is replaced by a strict inequality.
Finally, a (strictly) monotonic such set is homeo-monotonic with respect to $A^{\prime}$ if
$s_{i}(a_{i},a_{j})$ $\geq$ $s_{j}(a_{i},a_{j})$ for all $1\leq i<j\leq|V^{\prime}|$.
Verbally, when $k$ agents have identical preferences over their top $k$ items, and agents can be (strictly) ordered in how much they prefer any two consecutively-ranked items in this subset [cf (strict) monotonicity], then this ordering cannot be violated in how much they prefer any non-consecutively ranked items in this subset (cf homeo-monotonicity).
Definition 10.
A profile $S\in\mathcal{\widehat{S}}$ is homeo-monotonic if, whenever $V$ and $A$ can be partitioned into subsets $V^{1},\ldots,V^{m}$ and $A^{1},\ldots,A^{m}$ where $V^{1}$ is envy-free with respect to $A^{1}$ and $V^{i>1}$ is monotonically top-preference-coincident with respect to $A^{i>1}$, then $V^{i>1}$ is homeo-monotonic with respect to $A^{i>1}$.
Definition 11.
A profile $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic if it induces partitions of agents and items relative to which it is homeo-monotonic.
Example 2.
The table below describes an effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where $n=6$ and where the envy-free and top-preference-coincident sets that partition $V$ and $A$ are $(V^{1},A^{1})=(\{v_{1}\},\{a_{1}\})$, $(V^{2},A^{2})=(\{v_{2},v_{3},v_{4}\},\{a_{2},a_{3},a_{4}\})$, $(V^{3},A^{3})=(\{v_{5},v_{6}\},\{a_{5},a_{6}\})$:
$\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$ $\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{3}$}\hskip 20.0pt}$
1 2 3 4 5 6
$s_{i}(a,b)=$ $i=$ ${\color[rgb]{1,0,1}a_{1}}\succ_{1}\cdots$ ${\color[rgb]{1,0,0}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}}\succ_{i}\cdots$ ${\color[rgb]{0,0,1}a_{5}\succ_{i}a_{6}}\succ_{i}\cdots$
15 $(a_{1},a_{6})$ $(a_{2},a_{1})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{5},a_{1})$ $(a_{5},a_{3})$
14 $(a_{2},a_{6})$ $(a_{2},a_{6})$ $(a_{3},a_{5})$ $(a_{3},a_{6})$ $(a_{5},a_{4})$ $(a_{5},a_{1})$
13 $(a_{1},a_{5})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{3})$ $(a_{5},a_{4})$
12 $(a_{3},a_{6})$ $(a_{2},a_{4})$ $(a_{3},a_{6})$ $(a_{2},a_{1})$ $(a_{5},a_{2})$ $(a_{5},a_{2})$
11 $(a_{2},a_{5})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$ $(a_{3},a_{1})$ $(a_{6},a_{1})$ $(a_{6},a_{3})$
10 $(a_{1},a_{4})$ $(a_{3},a_{6})$ $(a_{3},a_{1})$ $(a_{4},a_{1})$ $(a_{6},a_{4})$ $(a_{6},a_{1})$
9 $(a_{4},a_{6})$ $(a_{3},a_{5})$ $(a_{4},a_{5})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$
8 $(a_{3},a_{5})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{6},a_{3})$ $(a_{6},a_{4})$
7 $(a_{2},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$
6 $(a_{1},a_{3})$ $(a_{4},a_{1})$ $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{2},a_{4})$ $(a_{4},a_{3})$
5 $(a_{5},a_{6})$ $(a_{4},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{1})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$
4 $(a_{4},a_{5})$ $(a_{5},a_{1})$ $(a_{1},a_{5})$ $(a_{1},a_{6})$ $(a_{6},a_{2})$ $(a_{6},a_{2})$
3 $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{4},a_{1})$ $(a_{2},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$
2 $(a_{2},a_{3})$ $(a_{5},a_{6})$ $(a_{1},a_{6})$ $(a_{2},a_{3})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$
1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
$\blacklozenge$
Remark 1.
If $V$ is envy-free with respect to $A$, then $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
Clearly, any set of agents can be partitioned into groups according to whether distinct or identical alternatives are the most preferred within each group. The additional structure imposed by effectively homeo-monotonic profiles amounts to requiring that: (i) the preferences of agents with the same top-ranked item also coincide in how they rank their second-, third-, $\ldots$, $k$th-most preferred item for some $k\leq n$ that coincides with the number of agents in that group; (ii) the preferences of distinct such groups are sufficiently heterogeneous. While this condition is restrictive and one can easily construct profiles where intensity-efficient allocations exist but the condition is violated, the proof of the next result shows that it is helpful toward clarifying some important mechanisms for the existence of such allocations in the general case. Moreover, a weaker version of this condition that we consider later turns out to be necessary for such existence.
Theorem 2.
Every effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation.
Proof.
Consider a profile $S\in\mathcal{\widehat{S}}$ with the postulated structure. Let $V^{1}:=\{v^{1}_{1},\ldots,v^{1}_{|V^{1}|}\}$ be the envy-free set with respect to $A^{1}:=\{a^{1}_{1},\ldots,a^{1}_{|A^{1}|}\}$, and let $V^{i>1}:=\{i_{1},\ldots,i_{|V^{i}|}\}$ be the $i$th top-preference-coincident set with respect to $A^{i>1}:=\{i_{1},\ldots,i_{|A^{i}|}\}$, $i=2,\ldots,m$. By the profile assumption, $|A^{i}|=|V^{i}|$ for all $i\leq m$, and the collections $(V^{i})_{i=1}^{m}$ and $(A^{i})_{i=1}^{m}$ partition the sets $V$ and $A$, respectively. Without loss of generality we may assume that $a^{1}_{j}$ is the top-ranked item of agent $v^{1}_{j}$, $j=1,\ldots,|A^{1}|$ (for the general case we apply a permutation over $\{1,\ldots,|A^{1}|\}$ on the $a^{1}$-subscripts). Recalling our earlier notation, this amounts to writing $a_{1}^{1}\equiv b_{1}^{1}$, $a_{2}^{1}\equiv b_{2}^{1}$, etc. Also without loss, for $A^{i>1}$, $V^{i>1}$ we may assume that $a^{i}_{1}\succ_{j}a^{i}_{2}\succ_{j}\cdots\succ_{j}a^{i}_{|A^{i}|}$ for all $v_{j}\in V^{i>1}$. (Note: the special cases where $V^{1}=\emptyset$, $V^{1}=V$ or $V^{i}=V$ for some $i>1$ are allowed by the postulated $S\in\mathcal{\widehat{S}}$, and so will they be in the proof.)
Lemma 3.
If $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic, then an allocation $x$ is Pareto-efficient with respect to $P_{S}$ if and only if it satisfies the following:
  1. (i)
    $v_{j}\in V^{1}$ $\Rightarrow$ $x_{j}=b_{j}^{1}$.
  2. (ii)
    $v_{j}\in V^{i>1}$ $\Rightarrow$ $x_{j}\in A^{i>1}$.
Proof.
For the “if” part, let $x$ be an arbitrary allocation that satisfies (i) and (ii). Assume to the contrary that $x$ is not Pareto efficient. Let $y$ be another allocation that Pareto-dominates $x$. This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$. Let $V^{*}\subsetneq V$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$. By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$. Suppose $V^{*}=\emptyset$. From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$. If the envy-free set $V^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in V^{1}$. If $V^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $V^{i>1}$ is non-empty. Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in V^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction. Now suppose $V^{*}\neq\emptyset$. Since $y$ Pareto-dominates $x$, the set $V^{**}:=V\setminus V^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty. From the preceding argument we know that for every $v_{l}\in V^{**}$ it must be that $v_{l}\not\in V^{1}$ and $v_{l}\neq v_{1}^{i}\in V^{i>1}$, for all $i>1$. By $s_{l}(y_{l},x_{l})>0$ and $|A|=|V|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$. By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$. Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in V^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $V^{i>1}$ and $V^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$. By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$. Since, by the definition of top-preference-coincident sets, $s_{l}(a,a^{\prime})>0$ and $s_{k}(b,b^{\prime})>0$ holds for all $a\in A^{i>1},a^{\prime}\in A\setminus A^{i>1}$ and $b\in A^{j>1},b^{\prime}\in A\setminus A^{j>1}$, and since, by (ii), it holds that $x_{l}\in A^{i>1}$ and $x_{k}\equiv y_{l}\in A^{j>1}$, the postulate $s_{l}(y_{l},x_{l})>0$ leads to a contradiction. Therefore, $x$ is Pareto-efficient.
For the “only if” part, consider an arbitrary Pareto-efficient allocation $x$. Suppose to the contrary that (i) is false. That is, $x_{j}\neq b_{j}^{1}$ for some $v_{j}\in V^{1}$. Let $b_{j}^{1}=x_{l}$ for the relevant distinct agent $v_{l}\in V$ (such an agent exists because $|V|=|A|$). Either $v_{l}\in V^{1}$ or $v_{l}\in V^{i>1}$. In the former case, swapping $x_{j}$ and $x_{l}$ between the two agents Pareto-improves upon $x$, contradicting its postulated Pareto optimality. In the latter case, from the assumption that $V^{i>1}$ is top-preference-coincident with respect to $A^{i>1}$, and that $S$ is homeo-monotonic, we have $a^{i}_{t}\succ_{l}x_{l}\equiv b^{1}_{j}$ for every $a^{i}_{t}\in A^{i>1}$. Furthermore, in this case (again because $|V|=|A|$) there must also exist another agent $k\in V^{p}$, $p\neq i$, such that $x_{k}\in A^{i>1}$. There are two possibilities: $p=1$ or $p>1$. If $p=1$, then the above contradiction is obtained now for $x_{l}$ and $x_{k}$. If $p>1$, then, by homeo-monotonicity of $S$ and the implied structure of $(V^{l>1},A^{l>1})$, $(V^{p>1},A^{p>1})$, we have $s_{l}(x_{k},x_{l})>0$ and $s_{k}(x_{l},x_{k})>0$, which also contradicts the postulated Pareto-efficiency of $x$. The argument establishing necessity of (ii) is analogous and omitted. ∎
Next, consider a top-preference-coincident set $V^{i>1}$ with respect to $A^{i>1}$. By effective homeo-monotonicity, either such a non-empty set exists or the envy-free set $V^{1}$ coincides with $V$. In the latter case, the allocation $x^{*}\in\mathcal{A}$ defined by $x^{*}_{i}=b^{1}_{i}$ is obviously the unique Pareto-efficient allocation at $S$ and therefore, trivially, intensity-efficient. So, in the sequel we assume that $V^{i>1}\neq\emptyset$.
Denote by $a_{1}^{i}\succ_{i}a_{2}^{i}\succ_{i}\cdots\succ_{i}a_{|A^{i}|}^{i}$ the induced preference order on $A^{i>1}$ which, by the definition of $(V^{i>1},A^{i>1})$, is common across all agents in $V^{i>1}$. Define the binary relation $\trianglerighteq^{i}$ on $V^{i>1}$ by
$v_{j}\trianglerighteq^{i}v_{k}$ $\Longleftrightarrow$ $s_{j}(a^{i}_{l},a^{i}_{l+1})\geq s_{k}(a^{i}_{l},a^{i}_{l+1})\text{ for all }l=1,\ldots,|A^{i>1}|-1$
By the postulated monotonicity of $(V^{i>1},A^{i>1})$, $\trianglerighteq^{i}$ is a weak order on $V^{i>1}$. Let $\trianglerighteq_{i}^{*}$ be some extension of $\trianglerighteq^{i}$ into a linear order over $V^{i>1}$, and denote by $\trianglerighteq^{*i}(j)$ the $\trianglerighteq^{*i}$-rank of agent $v_{j}\in V^{i>1}$. By construction, $\trianglerighteq^{*i}$ coincides with $\trianglerighteq^{i}$ except where it breaks any equivalence ties that the latter relation may feature.
Now define $x^{*}\in\mathcal{A}$ by
$x_{j}^{*1}:=b_{j}^{1},\;j=1,\ldots,|V^{1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{1}$
$x_{j}^{*i>1}:=a^{i}_{l},\;j=1,\ldots,|V^{i>1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{i>1}\text{ and }\trianglerighteq^{*i}(j)=l$
That is, each agent in the set $V^{1}$ (which could be empty) is assigned their most preferred item, while the $\trianglerighteq^{*i}$-$l$th-ranked agent in $V^{i>1}$ is assigned their $l$th most preferred item. Recalling that $V^{1}$ is envy-free and each $V^{i>1}$ is top-preference-coincident, it follows from Lemma 3 that $x^{*}$ is Pareto-efficient. We will prove that it is also intensity-efficient.
To this end, suppose to the contrary that some Pareto-efficient allocation $y\in\mathcal{A}$ intensity-dominates $x^{*}$. Then, there are agents $v_{i},v_{j}\in V$ such that $(x_{i}^{*},x_{j}^{*})=(y_{j},y_{i})$ and $s_{j}(y_{j},y_{i})\equiv s_{j}(x_{i}^{*},x_{j}^{*})>s_{i}(x^{*}_{i},x_{j}^{*})>0$. By construction of $x^{*}$, $v_{i},v_{j}$ cannot belong to the same $V^{l\geq 1}$. For $l=1$ this is obvious because $x^{*i}=b^{1}_{i}$ and $x^{*j}=b^{1}_{j}$, contradicting the above postulate. For $l>1$, moreover, one has $x^{*}_{i},x_{j}^{*}\in A^{l>1}$ while items in $A^{l>1}$ are assigned to agents in $V^{l>1}$ according to the linear order $\trianglerighteq^{*l}$, which, by homeo-monotonicity of $S$, ensures that $s_{i}(x_{i}^{*},x_{j}^{*})\geq s_{j}(x_{i}^{*},x_{j}^{*})$, thereby contradicting that postulate. The remaining possibility is $v_{i}\in V^{l>1}$ and $v_{j}\in V^{k>1}$ for some $k\neq l$. By construction of $x^{*}$, however, we have $x_{i}^{*}\in A^{l>1}$ and $x_{j}^{*}\in A^{k>1}$ which, because $V^{l>1}$, $V^{k>1}$ are top-preference-coincident with respect to $A^{l>1}$, $A^{k>1}$ and $A^{l>1}\cap A^{k>1}=\emptyset$, implies $s_{i}(x_{i}^{*},x_{j}^{*})\equiv s_{i}(y_{j},y_{i})>0$ and $s_{j}(x_{i}^{*},x_{j}^{*})\equiv s_{j}(y_{j},y_{i})<0$. Thus, the postulate that $y$ intensity-dominates $x^{*}$ is contradicted in all situations. ∎
Example 2 (continued).
By Lemma 3, there are 12 Pareto-efficient allocations in this profile, namely those which assign $a_{1}$ to $v_{1}$ and alternate items $a_{2},a_{3},a_{4}$ among agents $v_{2},v_{3},v_{4}$ and $a_{5},a_{6}$ among $v_{5},v_{6}$. Out of these 12 allocations, only the following 2 are the (in fact, equivalent) intensity-efficient ones: $(a_{1},a_{2},a_{3},a_{4},a_{5},a_{6})$; $(a_{1},a_{2},a_{3},a_{4},a_{6},a_{5})$. $\blacklozenge$
The way in which the intensity-efficient allocation $x^{*}$ is constructed in the proof of Theorem 2 readily suggests the following stronger implication for a class of special cases:
Corollary 2.
An intensity-dominant—up to equivalence—allocation exists in any effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where every top-preference-coincident set has two agents.
Importantly, although effective homeo-monotonicity is a sufficient but not necessary condition for the existence of intensity-efficient allocations, we show next that the restriction it imposes within monotonic top-preference-coincident sets—if the profile induces such sets—is also necessary whenever these sets are strictly monotonic.
Theorem 3.
If a profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation, then every strictly monotonic top-preference-coincident set in it is homeo-monotonic.
Proof.
[…]
Lemma 4.
[…]
Proof.
The only difference between allocations $x$ and $x^{\prime}$ lies in how they permute items in $A^{\prime}$ to agents in $V^{\prime}$. Denote their restrictions on these subsets by $\widehat{x}$ and $\widehat{x}^{\prime}$, respectively. By Lemma 3 and the postulated intensity-efficiency (hence Pareto-efficiency) of $x$ it follows that $x^{\prime}$ is Pareto-efficient too. Since $x$ is intensity-efficient, it is not intensity-dominated by $x^{\prime}$. Hence, there is no pair $(a^{\prime},a^{\prime\prime})\in A^{\prime}\times A^{\prime}$ such that $(a^{\prime},a^{\prime\prime})=(\widehat{x}_{i},\widehat{x}_{j})=(\widehat{x}^{\prime}_{j},\widehat{x}^{\prime}_{i})$ for some $v^{\prime}_{i},v^{\prime}_{j}\in V^{\prime}$, making $x$ and $x^{\prime}$ incomparable by intensity-dominance. The argument in the proof of Theorem 2 following Lemma 3 establishes that there is no other Pareto-efficient allocation that intensity-dominates $x^{\prime}$ either. Thus, $x^{\prime}$ is intensity-efficient. ∎
[…]
$\begin{array}[]{lllllllll}s^{\prime}_{1}(a^{\prime}_{i},a^{\prime}_{i+1})&>&s^{\prime}_{2}(a^{\prime}_{i},a^{\prime}_{i+1})&>&\cdots&>&s^{\prime}_{|V^{\prime}|}(a^{\prime}_{i},a^{\prime}_{i+1})&&\text{for all }i\leq|A^{\prime}|-1\end{array}$
Now recall that the allocation $x^{\prime}$ constructed above is intensity-efficient. This assigns $x^{\prime}_{m}=a^{\prime}_{m}\in A^{\prime}$ to each $v^{\prime}_{m}\in V^{\prime}$. Since $(V^{\prime},A^{\prime})$ is strictly monotonic but not homeo-monotonic, there are $v^{\prime}_{j},v^{\prime}_{l}\in V^{\prime}$ such that $v^{\prime}_{j}\trianglerighteq^{\prime}v^{\prime}_{l}$, jl>1j-l>1 and $s^{\prime}_{l}(a^{\prime}_{j},a^{\prime}_{l})>s^{\prime}_{j}(a^{\prime}_{j},a^{\prime}_{l})$. Since $(x^{\prime}_{j},x^{\prime}_{l})=(a^{\prime}_{j},a^{\prime}_{l})$, this implies that the allocation $x^{\prime\prime}$ which is identical to $x^{\prime}$ except that $(x^{\prime\prime}_{j},x^{\prime\prime}_{l})=(a^{\prime}_{l},a^{\prime}_{j})$ intensity-dominates $x^{\prime}$, contradicting the postulated intensity-efficiency of the latter allocation. ∎
The next example explains why “strictly monotonic” cannot be replaced by “monotonic” in the statement of Theorem 3.
Example 3.
For $n=4$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as follows:
$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$
1 2 3 4
$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}$ $a_{4}\succ_{4}\cdots\succ_{4}a_{1}$
6 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{4},a_{1})$
5 $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{1},a_{3})$
4 $(a_{1},a_{3})$ $(a_{1},a_{3})$ $(a_{2},a_{4})$
3 $(a_{1},a_{2})$ $(a_{1},a_{2})$ $(a_{3},a_{4})$
2 $(a_{2},a_{3})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
1 $(a_{3},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$
The Pareto-efficient allocations here are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$. Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$ and $z$ (these are dominated by $p$ and $y$, respectively). Yet $V^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$. $\blacklozenge$
Our last example, finally, shows that a strictly monotonic profile that is not homeo-monotonic may fail to have an intensity-efficient allocation.
Example 4.
For $n=5$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4},v_{5}\},\{a_{4},a_{5}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\}$, $\{a_{1},a_{2},a_{3}\})$, as follows:
$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{1}$\hskip 20.0pt}}$
1 2 3    4    5
$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}\succ_{i}a_{5}$ $a_{4}\succ_{4}\cdots$ $|$ $a_{5}\succ_{5}\cdots$
10 $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{4},\cdot)$ $(a_{5},\cdot)$
9 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $\vdots$ $\vdots$
8 $(a_{2},a_{5})$ $(a_{2},a_{5})$ $(a_{1},a_{3})$
7 $(a_{2},a_{4})$ $(a_{1},a_{3})$ $(a_{2},a_{5})$
6 $(a_{1},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$
5 $(a_{3},a_{5})$ $(a_{3},a_{5})$ $(a_{3},a_{5})$
4 $(a_{1},a_{2})$ $(a_{3},a_{4})$ $(a_{3},a_{4})$
3 $(a_{2},a_{3})$ $(a_{1},a_{2})$ $(a_{4},a_{5})$
2 $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
1 $(a_{4},a_{5})$ $(a_{4},a_{5})$ $(a_{2},a_{3})$
The Pareto-efficient allocations here are
$\begin{array}[]{lllllll}\underbrace{(c,a,b,d,e)}_{o},&\underbrace{(c,b,a,d,e)}_{t},&\underbrace{(a,b,c,d,e)}_{x},&\underbrace{(a,c,b,d,e)}_{y},&\underbrace{(b,c,a,d,e)}_{w},&\underbrace{(b,a,c,d,e)}_{z}\end{array}$
Yet the intensity-dominance relation $D$ is cyclic over this set:
[…]
$\blacklozenge$
Together, Theorems 2-3 and Examples 2-4 highlight the importance of the intuitive structure imposed by homeo-monotonicity for the general existence of intensity-efficient allocations, both as a sufficient condition and, in a weaker sense, as a necessary one too. As the counterexample for $n=4$ in the proof of Theorem 1 shows, moreover, a profile may fail to have an intensity-efficient allocation even when homeo-monotonicity as defined above is satisfied, yet a subtle variation of it is not. More specifically, that profile features: (i) $n=4$; (ii) $a\succ_{i}b\succ_{i}c\succ_{i}d$ for $i\leq 4$ (hence a top-preference-coincident set $V$ with respect to $A$); (iii) $s_{1}=s_{4}$, $s_{2}=s_{3}$. Although the four agents’ intensities over their common preferences do not define a monotonic top-preference-coincident set, one observes that emergent here is a variant of this notion whereby each set in the partition $\{\{v_{1},v_{2}\},\{v_{3},v_{4}\}\}$ of $V$ is monotonically top-preference-coincident with respect to $\{a,b,c\}\subset A$: $s_{1}(a,b)>s_{2}(a,b)$, $s_{1}(b,c)=s_{2}(b,c)$ and $s_{4}(a,b)>s_{3}(a,b)$, $s_{4}(b,c)=s_{3}(b,c)$. However, both these pairs of agents violate what homeo-monotonicity would require here: $s_{1}(a,c)<s_{2}(a,c)$ and $s_{4}(a,c)<s_{3}(a,c)$. Thus, while the problem of fully characterizing the conditions under which strict profiles have intensity-efficient allocations remains open, the preceding analysis hopefully provides non-trivial and useful guidance in that direction.
Figure 1: Simulations-based histograms showing the proportions of Pareto-efficient allocations that are also intensity-efficient when the number of agents and items is small.
Additional motivation for further study is also contributed by the analysis of uniform-randomly drawn intensity profiles when $n=3,4,5$. Figure 1 summarizes this output and displays how the distribution of the per-profile proportion of Pareto-efficient allocations that are also intensity-efficient varies with $n$. Interestingly, while the average number of Pareto-efficient allocations increases rapidly from 2.8 to 5.9 to 14.1 when $n=3,4,5$, respectively, the corresponding averages for intensity-efficient allocations are 1.6, 1.9 and 2.4. The refinement gains afforded by intensity-efficiency are therefore significant and appear to be increasing in $n$, with the mean proportion of intensity-efficient out of all Pareto-efficient allocations falling from 0.63 ($n=3$) to 0.2 ($n=5$). Finally, non-existence of intensity-efficient allocations is a rare occurrence in these simulations, with only 1 and 15 out of the respective 100,000 randomly drawn profiles being associated with a cyclic intensity-dominance relation over Pareto-efficient allocations when $n=4$ and $n=5$, respectively.
We end this section by remarking that the paper’s new analytical environment necessitated novel arguments in the proofs of this section’s main results. More specifically, viewed as an abstract mathematical problem, existence of intensity-efficient allocations amounts to identifying sufficient conditions for acyclicity of the intensity-dominance binary relation defined on the set of permutations over alternatives. The mathematical literature on the combinatorics of permutations has studied partial orders over such sets, which may even have the richer structure of a lattice that additionally ensures the existence of both a greatest (hence maximal) and a smallest element. These results, however, are not applicable in our environment. The reason is that those partial orders, namely the strong and weak Bruhat orders (Stanley, 2012, pp. 399-400), relate two permutations if and only if there is a single transposition (“flip”) between elements in the two permutations, and this transposition corresponds to an inversion (Stanley, 2012, p. 30) relative to some postulated fixed linear order over its elements. The weak Bruhat order (see Ceballos and Pons (2024) for a recent generalization) gives rise to a lattice—whose graph is known as the permut(a)(o)hedron—but under the additional assumption that the said single transposition occurs between adjacent elements of the two relevant permutations. Our problem on the other hand does not assume a fixed linear order over items and is defined by intensity-dominance comparisons that generally feature possibly multiple transpositions of non-adjacent elements, which are defined by dominance between specific entries in the two relevant columns of an integer matrix, defined in turn by the corresponding agents’ intensity orderings.
5 Eliciting Ordinal Intensities and Finding Intensity-Dominating Allocations
We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems. In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency in the intensity-dominating direction. Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational. To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
Proposition 3.
Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in exactly $\frac{1}{2}(n-1)(n-2)$ steps, hence in $O(n^{2})$ time.
Proof.
Define $f(n):=\frac{1}{2}(n-1)(n-2)$. Observe that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains
$|\dot{\succsim}_{l}|$ $=$ $2{n\choose 2}+n$ (31)
comparisons between pairs in $A\times A$. We proceed by induction.
Let $n=3$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$. By (31), $|\dot{\succsim}_{l}|=9$. By (6) and the assumption that $\succ_{l}$ is induced by $\dot{\succsim}_{l}$ we readily get $(a_{1},a_{3})\dot{\succsim}_{l}(a_{1},a_{2})$ and $(a_{1},a_{3})\dot{\succsim}_{l}(a_{2},a_{3})$. By (5), moreover, we have $(a_{1},a_{1})\dot{\sim}_{l}(a_{2},a_{2})\dot{\sim}_{l}(a_{3},a_{3})$. By (5), $(\mathbf{5}_{D})$ and $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$ we also get $(a_{1},a_{2})\dot{\succ}_{l}(a_{i},a_{i})$ and $(a_{2},a_{3})\dot{\succ}_{l}(a_{i},a_{i})$, $i\leq 3$. By (5) and the above comparisons, finally, we also have $(a_{i},a_{i})\dot{\succ}_{l}(a_{2},a_{1})\dot{\succ}_{l}(a_{3},a_{1})$ and $(a_{i},a_{i})\dot{\succ}_{l}(a_{3},a_{2})\dot{\succ}_{l}(a_{3},a_{1})$. Together, these account for 7 out of the 9 comparisons in $\dot{\succsim}_{l}$, with those between $(a_{1},a_{2})$, $(a_{2},a_{3})$ and between $(a_{3},a_{2})$, $(a_{2},a_{1})$ being the two missing ones. By (5) once again, knowing either the first or the second such comparison also pins down the other. Hence, $f(3)=1$ is indeed the number of steps needed to elicit $\dot{\succsim}_{l}$ from $\succ_{l}$ when $n=3$.
Now let $n=4$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}\succ_{l}a_{4}$. By the preceding arguments we know that, out of the $|\dot{\succsim}_{l}|=16$ total comparisons here, missing are the following 3: $(a_{1},a_{2})$ vs $(a_{2},a_{3})$; $(a_{2},a_{3})$ vs $(a_{3},a_{4})$; $(a_{1},a_{3})$ vs $(a_{2},a_{4})$. Hence, $f(4)=3$ is the number of necessary steps here too.
For the induction step of the proof when $n>4$, suppose $f(m)$ is the correct number for every $n\leq m$. We will show that $f(m+1)$ is the correct number when $n=m+1$. To this end, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m}$ and notice that, from the induction hypothesis and the information derived from the $n=3$ and $n=4$ cases above, we can decompose the $f(m)$ comparisons as follows:
(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-2$ (m2m-2 comparisons);
(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-3$ (m3m-3 comparisons);
(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-4$ (m4m-4 comparisons);
$\cdots$
(m2m-2):  $(a_{i},a_{i+m-1})$ vs $(a_{i+1},a_{i+m})$ (1 comparison).
Indeed, summing the terms inside brackets across the right hand sides of $(1)\cdots(m-2)$ we get
$\sum\limits_{i=2}^{m-1}(m-i)$ $=$ $\sum\limits_{i=1}^{m-1}(m-i)-(m-1)$
$=$ $\frac{1}{2}(m-1)m-(m-1)$
$=$ $(m-1)\left(\frac{1}{2}m-1\right)$
$=$ $\frac{1}{2}(m-1)(m-2)$
$=$ $f(m)$
Consider now $n=m+1$ and, once again without loss, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m+1}$. From the preceding decomposition of the required comparisons when $n=m$ we can deduce that, for $n=m+1$,
(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-1$ (m1m-1 comparisons);
(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-2$ (m2m-2 comparisons);
(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-3$ (m3m-3 comparisons);
$\cdots$
(m1m-1):  $(a_{i},a_{i+m})$ vs $(a_{i+1},a_{i+m+1})$ (1 comparison).
Summing across categories now yields $\sum\limits_{i=1}^{m}(m-i)=\frac{1}{2}(m-1)m=f(m+1)$. ∎
Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available. The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm. The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run. TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment. Shapley and Scarf (1974) proved that TTC leads to a core allocation. Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique. Extending the results in Roth (1982), Ma (1994) later established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes. This mechanism is not strategy-proof. Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism. The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available. These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority under some ranking over agents.
Bogomolnaia and Moulin (2001), finally, demonstrated that, despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, these are generally ex ante inefficient in the sense that, based on the stochastic-dominance relation defined by all agents’ preferences, they may be dominated by another random allocation. In response to this realization, the authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
Informed by the main insights of the above-cited literature, we now consider the following algorithm for the problem at hand:
Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
Step 1: Choose a linear order $\gg$ over $V$ from the uniform distribution with support $n!$.
Step 2: Apply RP under $\gg$ on $P_{S}$ to arrive at a Pareto-efficient allocation $x$.
Step 3: Reorder the entries in $x$ according to $\gg$, with $j>i$ for $x_{i},x_{j}$ if and only if $i\gg j$.
Step 4: Search for the $\gg$-lowest agent $v_{j}\in\{v_{2},\ldots,v_{n}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$. If such an agent exists, swap $x_{1}\leftrightarrow x_{j}$ and remove $v_{1},v_{j}$ from $V$ and $x_{1},x_{j}$ from $A$.
If no such agent exists, remove $v_{1}$, $x_{1}$.
Repeat Step 4 until no agents remain.
Proposition 4.
If $S\in\mathcal{\widehat{S}}$ and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm terminates in $O(n^{2})$ time and either produces the Pareto-efficient allocation that it finds in its RP-implementing Step 2 or an allocation that intensity-dominates it.
Proof.
Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step. In the worst-case scenario, Step 4 always—with the possible exception of the agent ranked second-last by $\gg$—performs exhaustive search and removes one agent and one object in every iteration: Starting with $v_{1},x_{1}$, after n1n-1 searches it removes $v_{1},x_{1}$. Continuing with $v_{2},x_{2}$, after n2n-2 searches it removes $v_{2},x_{2}$. $\cdots$ Continuing with $v_{n-1},x_{n-1}$, after one search it removes both $v_{n-1},x_{n-1}$ and $v_{n},x_{n}$, either by implementing the swap $x_{n-1}\leftrightarrow x_{n}$ or otherwise. Worst-case performance is therefore associated with a total number of searches equalling
$\sum_{m=1}^{n-1}(n-m)$ $=$ $n^{2}-n-\sum_{m=1}^{n-1}m$
$=$ $n^{2}-\sum_{m=1}^{n}m$
Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in $O(n^{2})$ time.
It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$. Suppose $n=3$ first. Write $x:=(x_{1},x_{2},x_{3})$. By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap. Thus, $y$ must be one of the following:
(i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$
(ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$
(iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$
In each case, $yDx$. Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic. Denote by $(V^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $V$ and $A$ into a (possibly empty) top-preference diverse set $V^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $V^{i>1}$ with respect to $A^{i>1}$, $i\leq k$. By the properties of these sets profiles, and by the definition of algorithm’s Step 4, irrespective of the $\gg$-order on $V$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in V^{1}$. Thus, $y_{i}=x_{i}$ for all $v_{i}\in V^{1}$. By the properties of $(V^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in V^{l>1}$ and $v_{j}\in V^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct. It follows that $y$ differs from $x$ by at most $\frac{n}{2}$ or $\frac{n-1}{2}$ pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such swap $x_{i}\leftrightarrow x_{j}$ occurring within the same top-preference-coincident pair $(V^{l>1},A^{l>1})$. By Step 4, finally, every such swap reassigns the jointly preferred (by Pareto-efficiency of $x$) item within the relevant pair in the direction of intensity-dominance. Hence, $yDx$. ∎
The allocation delivered by this algorithm is not intensity-efficient in general. This is so because the forced exit of agents and items in the fourth step—a crucial feature for the algorithm’s manageable quadratic-time complexity—means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation. Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP. It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.
In light of the discussion preceding the statement of Proposition 4, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency. Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 3, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully. The answer is negative. To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$. The possible orderings over $V$ are: 1. $v_{1}\gg v_{2}\gg v_{3}$; 2. $v_{1}\gg v_{3}\gg v_{2}$; 3. $v_{2}\gg v_{1}\gg v_{3}$; 4. $v_{2}\gg v_{3}\gg v_{1}$; 5. $v_{3}\gg v_{1}\gg v_{2}$; 6. $v_{3}\gg v_{2}\gg v_{1}$. The two Pareto-efficient allocations are $(a,b,c)$ and $(b,a,c)$, and the former intensity-dominates the latter. Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with probability $\frac{1}{2}$). If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it. However, if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e. with probability $\frac{1}{2}$) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem. Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up. Abdulkadiroğlu et al. (2011), however, followed by Pycia (2011); Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for. In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools. In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible. In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.